Find the unit vector of A 3i -8j +5k


1
Expert's answer
2021-08-02T08:39:37-0400
"A=\sqrt{3^2+8^2+5^2}=7\sqrt{2}\\
\vec{e}_A=\frac{1}{7\sqrt{2}}(\bold{3i -8j +5k})"